The probabilities that a customer buys 2,3,4,5 or 6 items in the convenience store a 0.32 0.12 0.23 0.18 and 0.15 respectively

1
Expert's answer
2022-03-09T11:58:12-0500

we define the following

mean= ( (2 * 0.32) + (3 * 0.12) + (4 * 0.23) + (5 * 0.18) + (6 * 0.15) ) = 3.72

variance= ( (22 * 0.32) + (32 * 0.12) + (42 * 0.23) + (52 * 0.18) +(62 * 0.15) ) - (3.722)

=15.94 - 13.8384 = 2.1016


standard deviation = (variance)1/2 = (2.1016)1/2 = 1.449689622